# V Holešovičkách 2, 18000 Praha 8, Czech Republic

### v1

@@ C.4 Sample contour plots
For eccentricities $e\gtrapprox 0.9497$ (or, equivalently, for ellipticities $\gtrapprox 0.687$), the changing structure of the contour plots can be used to identify all the additional parameter-space boundaries seen in the right panel of Fig. 8 and in Fig. 10. In the three columns of Fig. 20 we present sample contour plots for eccentricities $e\in\{0.98,0.9881,0.9986\}$, which are also marked for orientation by the top three horizontal dashed lines in the parameter-space maps in Fig. 8 and Fig. 10. The top row of Fig. 20 shows the tangential contour plots, the middle row shows the central part of the radial contour plots, and the bottom row shows the large-scale radial contour plots out to the respective radial limiting curves. The contours in the top row correspond to the same $\kappa_{\text{s}}$ values and spacings as in the top panel of Fig. 19, except in the right column, in which we included only integer-valued contours for $\kappa_{\text{s}}>1$ to avoid crowding. In the middle row, contours starting from $\kappa_{\text{s}}=0.5$ are plotted for the same values and spacings as in the bottom panel of Fig. 19, except in the right column, in which we included only contours from $\kappa_{\text{s}}=1$ to $\kappa_{\text{s}}=10$. In the bottom row, contours are plotted from $\kappa_{\text{s}}=1$ to $\kappa_{\text{s}}=20$, with bold black contours marking multiples of 5.
@@
The radial critical curves seen in the bottom two rows of Fig. 20 show less variability, with the same character of the central inflection and cusp curves as in the bottom panel of Fig. 19. What can be seen newly in the bottom and middle right plots are the orange radial limiting curves, defined by Eq. (92) and with dimensions plotted in Fig. 13. Note that the inflection curve has an outer component lying beyond the radial limiting curve in the left two columns, and thus of no relevance to the corresponding radial critical curves. However, for $e=0.9986$ in the right column this component of the inflection curve reaches inside the radial limiting curve, so that radial critical curves with $\kappa_{\text{s}}\gtrapprox 54.3$ are pinched in near the minor axis. The corresponding second radial “peanut” transition is indicated by the dashed purple boundary in Fig. 10. This highest-eccentricity feature of the ellipsoidal NFW lens occurs only for eccentricities $e\gtrapprox 0.9946$. At lower eccentricities this component of the inflection curve does not cross the radial limiting curve and radial critical curves remain fully convex for any $\kappa_{\text{s}}$ above the solid purple boundary in Fig. 8.

### v2

@@ C.4 Sample contour plots
For eccentricities $e\gtrapprox 0.9497$ (or, equivalently, for ellipticities $\gtrapprox 0.687$), the changing structure of the contour plots can be used to identify all the additional parameter-space boundaries seen in the right panel of Fig. 8 and in Fig. 20. In the three columns of Fig. 19 we present sample contour plots for eccentricities $e\in\{0.98,0.9881,0.9986\}$, which are also marked for orientation by the top three horizontal dashed lines in the parameter-space maps in Figs. 8 and 20. The top row of Fig. 19 shows the tangential contour plots, the middle row shows the central part of the radial contour plots, and the bottom row shows the large-scale radial contour plots out to the respective radial limiting curves. The contours in the top row correspond to the same $\kappa_{\text{s}}$ values and spacings as in the top panel of Fig. 18, except in the right column, in which we included only integer-valued contours for $\kappa_{\text{s}}>1$ to avoid crowding. In the middle row, contours starting from $\kappa_{\text{s}}=0.5$ are plotted for the same values and spacings as in the bottom panel of Fig. 18, except in the right column, in which we included only contours from $\kappa_{\text{s}}=1$ to $\kappa_{\text{s}}=10$. In the bottom row, contours are plotted from $\kappa_{\text{s}}=1$ to $\kappa_{\text{s}}=20$, with bold black contours marking multiples of 5.
@@
The radial critical curves seen in the bottom two rows of Fig. 19 show less variability, with the same character of the central inflection and cusp curves as in the bottom panel of Fig. 18. What can be seen newly in the bottom and middle right plots are the orange radial limiting curves, defined by Eq. (92) and with dimensions plotted in Fig. 12. We note that the inflection curve has an outer component lying beyond the radial limiting curve in the left two columns, and thus of no relevance to the corresponding radial critical curves. However, for $e=0.9986$ in the right column this component of the inflection curve reaches inside the radial limiting curve, so that radial critical curves with $\kappa_{\text{s}}\gtrapprox 54.3$ are pinched in near the minor axis. The corresponding second radial “peanut” transition is indicated by the dashed purple boundary in Fig. 20. This highest-eccentricity feature of the ellipsoidal NFW lens occurs only for eccentricities $e\gtrapprox 0.9946$. At lower eccentricities this component of the inflection curve does not cross the radial limiting curve and radial critical curves remain fully convex for any $\kappa_{\text{s}}$ above the solid purple boundary in Fig. 8.
@@ C.5 Final comments
Appendix D Lensing regimes in parameter space at high $\kappa_{\text{s}}$
The structure of the parameter-space map of the ellipsoidal NFW lens at higher convergence-parameter values, further to the right of the range of Fig. 8 is shown in Fig. 20 for $\kappa_{\text{s}}\in[0,140]$. We show here only the region of higher ellipticity, $1-\sqrt{1-e^{2}}\geq 0.65$, corresponding to eccentricities $e\gtrsim 0.937$. No boundaries occur at lower eccentricity for $\kappa_{\text{s}}>5$ within the range of Fig. 20. The plot illustrates the changes of the boundaries from Fig. 8 as well as the appearance of another new boundary. The thin dashed purple boundary extending to the right from $\kappa_{\text{s}}\approx 41.6$ corresponds to another “peanut” transition on the radial critical curve. In this one two symmetric pairs of inflection points appear at the minor axis of the mass distribution, so that to the right of the boundary the radial critical curve is pinched in vertically.
Regarding the sequence of the boundaries, a closer inspection reveals there are only two boundaries that gradually intersect others within the range of Fig. 20; the remaining boundaries do not cross each other. The solid black cusp-piercing boundary intersects the green dotted “bulge” boundary at $\kappa_{\text{s}}\approx 7.37$, the bold green dashed butterfly boundary at $\kappa_{\text{s}}\approx 26.1$, and the purple dashed radial “peanut” boundary at $\kappa_{\text{s}}\approx 121$. The solid purple radial “peanut” boundary crosses the thin green dashed “peanut” boundary at $\kappa_{\text{s}}\approx 2.98$, the bold green dashed butterfly boundary at $\kappa_{\text{s}}\approx 7.62$, and the dashed purple radial “peanut” boundary $\kappa_{\text{s}}\approx 78.3$. Over most of the range of Fig. 20 the dotted green “bulge” boundary represents the first transition when increasing the eccentricity from zero. The boundary drops to a minimum at $(\kappa_{\text{s}},e)\approx(62.6,0.950)$, then rises slowly with increasing $\kappa_{\text{s}}$. In general, we note the entirely different sequence of nine transitions with increasing eccentricity at the right edge of Fig. 20 in comparison with the generic low-convergence sequence of four transitions as seen in the left panel of Fig. 8.
We have not explored the parameter space for higher values of $\kappa_{\text{s}}$ beyond the range of Fig. 20 comprehensively. Overall, the boundaries tend to level out, their rate of change declines with increasing $\kappa_{\text{s}}$. The lower part of the new purple dashed peanut boundary from Fig. 20 declines monotonically asymptotically to $e\approx 0.9945$. The declining lower part of the bold green dashed butterfly boundary eventually crosses the slowly rising dotted green bulge boundary at $(\kappa_{\text{s}},e)\approx(8000,0.956)$. For higher values of $\kappa_{\text{s}}$ the butterfly boundary marks the first transition to occur when increasing the eccentricity from zero. However, at such high values of the convergence parameter these results purely illustrate the changing properties of the simple lens model.
Figure 20: High-$\kappa_{\text{s}}$ lensing-regime boundaries of the ellipsoidal NFW lens mapped in its parameter space, as described in Appendix D, plotted here for ellipticities $0.651-\sqrt{1-e^{2}}>$ 0.65 and for convergence parameters up to $\kappa_{\text{s}}=140$. Notation same as in Fig. 8, with the additional thin dashed purple boundary marking the minor-axis radial-peanut transition.